# Contextual Continuum Bandits: Static Versus Dynamic Regret

### v3

@@ Abstract
We study the contextual continuum bandits problem, where the learner sequentially receives a side information vector and has to choose an action in a convex set, minimizing a function associated with the context. The goal is to minimize all the underlying functions for the received contexts, leading to a dynamic (contextual) notion of regret, which is stronger than the standard static regret. Assuming that the objective functions are Hölder with respect to the contexts, we demonstrate that any algorithm achieving a sub-linear static regret can be extended to achieve a sub-linear dynamic regret. We further study the case of strongly convex and smooth functions when the observations are noisy. Inspired by the interior point method and employing self-concordant barriers, we propose an algorithm achieving a sub-linear dynamic regret. Lastly, we present a minimax lower bound, implying two key facts. First, no algorithm can achieve sub-linear dynamic regret over functions that are not continuous with respect to the context. Second, for strongly convex and smooth functions, the algorithm that we propose achieves, up to a logarithmic factor, the minimax optimal rate of dynamic regret as a function of the number of queries.
@@ 1 Introduction
$\mathbf{E}\Big{[}\sum_{t=1}^{T}f_{t}(\boldsymbol{z}_{t})-\min_{\boldsymbol{z}\in\Theta}\sum_{t=1}^{T}f_{t}(\boldsymbol{z})\Big{]}.$ (1)
However, in this paper we consider a more ambitious task, which is to minimize the dynamic regret
$\mathbf{E}\Big{[}\sum_{t=1}^{T}f_{t}(\boldsymbol{z}_{t})-\sum_{t=1}^{T}\min_{\boldsymbol{z}\in\Theta}f_{t}(\boldsymbol{z})\Big{]}.$ (2)
Note that (2) presents a stronger notion of regret; achieving sub-linear dynamic regret implies sub-linear static regret. The difference between static and dynamic regret is where the minimum is placed. If static regret is small, it means the algorithm is at least as good as a single constant action that does well on the average of the functions seen during learning. In contrast, proving that dynamic regret is small implies that the algorithm is good for any function individually.
There is a sharp difference between static and dynamic regret. In general, obtaining a sub-linear static regret does not imply convergence properties for the sequence $\{f_{t}(\boldsymbol{z}_{t})\}_{t=1}^{T}$. The static regret is not necessarily positive. Conversely, achieving sub-linear dynamic regret (2) implies that for any $\epsilon>0$, only for a finite number of rounds we have
$\mathbf{E}\big{[}f_{t}(\boldsymbol{z}_{t})-\min_{\boldsymbol{z}\in\Theta}f_{t}(\boldsymbol{z})\big{]}\geq\epsilon.$
In other words, if an algorithm achieves sub-linear dynamic regret, it asymptotically approaches the minimum values of the received functions. Therefore a key question arises: When it is possible to achieve a sub-linear dynamic regret?
In the setting of non-bandit online zero- and first-order optimization, several authors have extensively explored the minimization of dynamic regret (2), see e.g., [34, 54, 9, 19, 24, 52]. To achieve sub-linear dynamic regret, these works introduce complexity assumptions on the received functions. Such assumptions characterize similarity between two consecutive functions helping the learner to leverage previously received information for subsequent inference. In the present paper, we consider the problem of contextual continuum bandits under the similarity notion described by Assumption 1 below.
@@
  • The learner chooses a query point $\boldsymbol{z}_{t}\in\Theta$ and gets a noisy evaluation of $f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})$:
@@
$\sum_{t=1}^{T}\mathbf{E}\left[f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})-\text{min}_{\boldsymbol{z}\in\Theta}f(\boldsymbol{z},\boldsymbol{c}_{t})\right].$ (4)
By letting $f_{t}(\cdot)=f(\cdot,\boldsymbol{c}_{t})$ we see that (4) is equivalent to the dynamic regret (2). Here and in the sequel, unless other assumption is explicitly stated, we assume that $\boldsymbol{c}_{t}$ is drawn by an adversary who may be aware of all previous actions of the learner and past information received by the learner. This formalization is an extension of the contextual multi-armed bandit problem [5] to the continuous action setup. One can consider the continuous dosage drug design problem as an example. At each round $t$, the learner encounters a patient with a specific medical profile encoded as $\boldsymbol{c}_{t}$. The learner’s goal is to craft a drug containing $d$ substances that closely matches the optimal drug design for a patient with the profile $\boldsymbol{c}_{t}$. Informally, achieving a sub-linear dynamic regret implies that, as $t$ increases, the designed drug approaches the best possible drug for the medical profile $\boldsymbol{c}_{t}$. However, in Section 5 we demonstrate that, regardless of the behavior of the function $f(\cdot,\boldsymbol{c})$, if the function $f(\boldsymbol{x},\cdot)$ lacks continuity, obtaining a sub-linear dynamic regret may be impossible. Hence, we are lead to assume that $f(\boldsymbol{x},\cdot)$ is a continuous function for any $\boldsymbol{x}\in\mathbb{R}^{d}$. This is aligned with the example of continuous dosage drug design. Indeed, it is natural to consider that the effect of a fixed drug is similar for two patients with similar medical profiles. This intuition leads us to the following somewhat stronger assumption that we adopt in the sequel.
Assumption 1.
@@
In Section 3, we provide a static-to-dynamic regret conversion theorem that can be summarized as follows. Consider functions $f_{t}$ that can be represented as $f_{t}(\boldsymbol{x})=f(\boldsymbol{x},\boldsymbol{c}_{t})$, where $f:\mathbb{R}^{d}\times[0,1]^{p}\to\mathbb{R}$. We show that if there exists an algorithm attaining a sub-linear static regret (1) when all $f_{t}$’s belong to a class of functions $\mathcal{F}$, then the same algorithm can be extended to achieve sub-linear dynamic regret for functions within $\mathcal{F}\cap\mathcal{F}_{\gamma}(L)$.
In Section 4, we focus on a specific class $\mathcal{F}$ that consists of strongly convex and smooth functions. In order to apply the approach developed in Section 3, we first construct an algorithm that achieves sub-linear static regret, see Algorithm 2. It is inspired by the Bandit Convex Optimization (BCO) algorithm proposed in [20], which is valid for non-noisy observations, that is, when $\xi_{t}\equiv 0$ in (3). In contrast, Algorithm 2 works for noisy observations. Next, for functions $f$ that additionally belong to the class $\mathcal{F}_{\gamma}(L)$, we extend the algorithm to attain a sub-linear dynamic regret using the conversion theorem of Section 3. Finally, in Section 5 we demonstrate that, if $f(\boldsymbol{x},\cdot)$ lacks continuity, obtaining a sub-linear dynamic regret may be impossible. Moreover, the minimax lower bound established in Section 5 implies near optimality (up to a logarithmic factor) of the algorithm proposed in Section 4.
@@ 2 Related work
The material presented below can be approached from three interconnected points of view: (i) the setting of contextual bandits, which is the main focus of this paper; (ii) the study of dynamic regret and the conditions under which sub-linear dynamic regret is achievable; and (iii) the convex bandits setting with strongly convex objective functions (see Sections 4 and 5).
Contextual bandits. Contextual bandit problem is a variant of bandit problem, where at each round the learner is equipped with an additional side information on the subsequent objective function. The problem has been introduced by [46] in the context of multi-armed bandits. Improvements and different setups, but all within the framework of multi-armed bandits, have been studied by many authors, see, e.g., [18, 45, 7, 39, 40, 38, 28, 6, 25, 48]. In the framework of continuum bandits, [42] studied the problem when Assumption 1 holds with $\gamma=1$, along with an additional assumption that the objective functions are Lipschitz with respect to the action variable, and proposed a method achieving an exponentially growing with $p$ dynamic regret. In [14], the authors studied the problem when Assumption 1 is satisfied, the objective functions are Lipschitz and convex with respect to the action variable, and the objective functions attain their global minimum in the constraint set $\Theta$. They proposed an algorithm that achieves a sub-linear dynamic regret and proved its optimality by deriving a matching minimax lower bound.
In a framework close to the one that we consider in Sections 4 and 5, [33] addressed the problem of contextual bandits when the functions are strongly convex and smooth. To establish an upper bound for the regret, [33] assumed that, for each $\boldsymbol{c}\in[0,1]^{p}$, the global minimizer $\operatorname*{arg\,min}_{\boldsymbol{x}\in\mathbb{R}^{d}}f(\boldsymbol{x},\boldsymbol{c})$ is an interior point of the constraint set $\Theta$. While this assumption might be hard to grant in practice, it does simplify the analysis considerably. In the present paper, we drop this assumption and obtain in Section 4 an upper bound that improves upon the result of [33]. Furthermore, in Section 5 we present a matching lower bound demonstrating that the algorithm that we propose is optimal to within a logarithmic factor.
@@
Dynamic regret. In the context of online convex optimization, numerous authors have obtained bounds on the dynamic regret. In the first-order optimization setup, where the learner can access a noisy gradient of the function at any given point, dynamic regret was explored by [9, 24, 34, 47, 49, 52, 53, 44]. For zero-order optimization, the dynamic regret has been studied by [54, 51, 50, 47, 15]. Among these papers, only [47, 15, 51] have addressed the classical bandit framework, in which the learner is not allowed to query function evaluations outside of the constraint set $\Theta$. We may also note that all the papers on contextual bandits cited above are dealing with the contextual regret, which is equivalent to the dynamic regret if the contexts are adversarial as discussed after (4). However, it is quite common in the literature to assume that the contexts are i.i.d. rather than adversarial.
Convex bandits. Convex bandits represent a family of constrained convex online optimization problems where, at each round, the learner can only observe a single function evaluation, and the query point must belong to the constraint set $\Theta$. Static regret in convex bandits problems has been extensively studied in the literature, see e.g., [32, 37, 30] and the references therein. A particular instance dealing with strongly convex and smooth functions has been investigated in [20] and [23] in a non-stochastic setting. In [20], the authors proposed an algorithm leveraging an interior point method and a self-concordant function. The achieved regret is nearly optimal, ranging from $d\sqrt{T}$ to $d^{3/2}\sqrt{T}$ depending on the geometry of the constraint set. Assuming that all the minimizers of the objective functions are in the interior of the constraint set, [23] introduced an algorithm with a regret of the order $d\sqrt{T}$. Both [20] and [23] studied the problem with no noise ($\xi_{t}\equiv 0$). In Section 4, we extend the results of [20] to noisy function evaluation as in (3).
Regarding simply convex functions, the first studies date back to [16, 26], where the authors addressed the setting with convex and Lipschitz objectives and proposed algorithms that achieve a regret of the order $T^{3/4}$ up to a factor depending on $d$. These works assume that the observed function values are noise-free. Successive improvements [12, 21, 8, 29] showed that a regret of the order $\sqrt{T}$ (up to a factor depending on $d$) can be attained in the noise-free setting, the sharpest known rate being $d^{2.5}\sqrt{T}$ [29]. Noisy variant of the problem was studied only when the objective function does not vary over time and the noise is sub-Gaussian. This led to algorithms with a regret of the order $d^{16}\sqrt{T}$ [2] and $d^{4.5}\sqrt{T}$ [31]. For a detailed survey on convex bandits, we refer the reader to [30, Sec. 2].
3 Static-to-dynamic regret conversion

### v6

# A conversion theorem and minimax optimality for continuum contextual bandits

@@ Abstract
We study the contextual continuum bandits problem, where the learner sequentially receives a side information vector and has to choose an action in a convex set, minimizing a function associated to the context. The goal is to minimize all the underlying functions for the received contexts, leading to the contextual notion of regret, which is stronger than the standard static regret. Assuming that the objective functions are $\gamma$-Hölder with respect to the contexts, $0<\gamma\leq 1$, we demonstrate that any algorithm achieving a sub-linear static regret can be extended to achieve a sub-linear contextual regret. We prove a static-to-contextual regret conversion theorem that provides an upper bound for the contextual regret of the output algorithm as a function of the static regret of the input algorithm. We further study implications of this general result for three fundamental cases of dependency of the objective function on the action variable: (a) Lipschitz bandits, (b) convex bandits, (c) strongly convex and smooth bandits. For Lipschitz bandits and $\gamma=1$, combining our results with the lower bound of Slivkins (2014), we prove that the minimax optimal contextual regret for the noise-free adversarial setting scales as $T^{(p+d+1)/(p+d+2)}$ up to logarithmic factors, where $p$ and $d$ are the dimensions of the context and of the action spaces, respectively, and $T$ is the number of queries. Then we prove that, when the evaluations are noisy, the rate of contextual regret in $T$ for convex bandits is the same as for strongly convex and smooth bandits and scales as $T^{(p+\gamma)/(p+2\gamma)}$ up to logarithmic factors.
Lastly, we present a minimax lower bound, implying two key facts. First, obtaining a sub-linear contextual regret may be impossible over functions that are not continuous with respect to the context. Second, for convex bandits and strongly convex and smooth bandits, the algorithms that we propose achieve, up to a logarithmic factor, the minimax optimal rate of contextual regret as a function of the number of queries.
@@ 1 Introduction
$R_{T}:=\mathbf{E}\Big{[}\sum_{t=1}^{T}f_{t}(\boldsymbol{z}_{t})-\min_{\boldsymbol{z}\in\Theta}\sum_{t=1}^{T}f_{t}(\boldsymbol{z})\Big{]}.$ (1)
In the present paper, we assume that the learner has access to side information about the loss function, namely a context at each round. Leveraging this additional information, the learner is expected to minimize another type of regret referred to as the contextual regret. We consider the problem of contextual continuum bandits stated as follows.
@@
  • Based on the past values $\{(y_{k},\boldsymbol{z}_{k},\boldsymbol{c}_{k})\}_{k=1}^{t-1}$, the learner chooses a query point $\boldsymbol{z}_{t}\in\Theta$ and gets a noisy evaluation of $f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})$:
@@
$R_{T}^{\mathrm{cntx}}:=\mathbf{E}\Big{[}\sum_{t=1}^{T}f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})-\sum_{t=1}^{T}\min_{\boldsymbol{z}\in\Theta}f(\boldsymbol{z},\boldsymbol{c}_{t})\Bigg{]}.$ (3)
Some remarks about this setting are in order. As a general framework for the problem, we assume that the contexts are drawn sequentially from the hypercube $[0,1]^{p}$. This presentation is chosen to simplify the notation. However, all the upper bounds on the regret throughout the paper hold if the contexts are drawn from a non-empty compact subset of $[0,1]^{p}$. Moreover, at first sight, by introducing the notation $f_{t}(\boldsymbol{z})=f(\boldsymbol{z},\boldsymbol{c}_{t})$, the contextual continuum bandit problem as stated above might appear as a special case of the non-contextual continuum bandit problem with dynamic regret (CBDR). In the CBDR setting, the learner selects the query point $\boldsymbol{z}_{t}$ at time $t$ based on past observations $\{y_{k},\boldsymbol{z}_{k}\}_{k=0}^{t-1}$ to minimize the dynamic regret:
${R_{T}^{\mathrm{dyn}}:=}\mathbf{E}\Big{[}\sum_{t=1}^{T}f_{t}(\boldsymbol{z}_{t})-\sum_{t=1}^{T}\min_{\boldsymbol{z}\in\Theta}f_{t}(\boldsymbol{z})\Big{]}.$ (4)
With this notation, we have ${R_{T}^{\mathrm{cntx}}=R_{T}^{\mathrm{dyn}}}$. However, the two settings are not equivalent. Indeed, the contextual setting that we consider is “statistically easier" than CBDR since partial information about $f_{t}$ in the form of the context $\boldsymbol{c}_{t}$ is revealed to the learner before the action is chosen. This suggests that a better behavior of the regret is potentially achievable.
Another performance measure for continuum contextual bandits called the $L$-regret was introduced by Hazan and Megiddo (2007):
$R_{T}^{L}:=\mathbf{E}\left[\sum_{t=1}^{T}f_{t}(\boldsymbol{z}_{t})-\min_{\boldsymbol{z}(\cdot)\in X_{L}}\sum_{t=1}^{T}f_{t}(\boldsymbol{z}(\boldsymbol{c}_{t}))\right].$ (5)
Here, the minimum is taken over the class $X_{L}$ of all actions satisfying the Lipschitz condition with respect to the context, with given Lipschitz constant $L$. For $L=0$ the $L$-regret is equal to the static regret $R_{T}$.
Considering $f_{t}(\boldsymbol{z})=f(\boldsymbol{z},\boldsymbol{c}_{t})$ we have that $R_{T}^{\mathrm{cntx}}=R_{T}^{\mathrm{dyn}}\geq R_{T}^{L}\geq R_{T}$. The setting with $L$-regret, where $L>0$, is more challenging than the static regret setting as constant actions trivially satisfy the Lipschitz condition. In turn, controlling the contextual regret $R_{T}^{\mathrm{cntx}}$ is more challenging than the $L$-regret since it evaluates the risk against all possible choices of actions and not only against the Lipschitz actions. By definition, small static regret indicates that the policy is at least as effective as a single constant action performing well on the average of all observed functions during learning. In contrast, small dynamic regret signifies that the policy performs well for each function individually. For non-contextual bandits, the CBDR problem has been recently investigated by Yang et al. (2016); Chen and Giannakis (2018); Zhao et al. (2021); Wei and Luo (2021); Wang (2023) under restrictions on various characteristics of the evolution of $f_{t}$’s over time, which are not interpretable in terms of the smoothness assumptions on $f(\cdot,\cdot)$ used in continuum contextual settings.
In Section 5, we demonstrate that, regardless of the behavior of the function $f(\cdot,\boldsymbol{c})$, if the function $f(\boldsymbol{x},\cdot)$ lacks continuity, obtaining a sub-linear contextual regret may be impossible. Hence, we focus on the setting where $f(\boldsymbol{x},\cdot)$ is a continuous function for any $\boldsymbol{x}\in\mathbb{R}^{d}$. Specifically, in the sequel we adopt the following slightly stronger assumption.
Assumption 1
@@
In this paper, we develop a meta-algorithm approach to continuum contextual bandit problem. In short it can be described as follows. Let $\mathcal{F}^{\prime}$ be a class of real-valued functions on $\mathbb{R}^{d}$. Assume that we are given an input algorithm $\pi$ achieving a sub-linear static regret in adversarial bandit setting over $T$ runs for the sequence of functions $\{f_{t}(\cdot)\}_{t=1}^{T}$, where each $f_{t}\in\mathcal{F}^{\prime}$. Then, we associate to it an output contextual bandit algorithm with sub-linear contextual regret for any loss $f(\cdot,\cdot)$ belonging to the class $\mathcal{G}_{\gamma}(\mathcal{F}^{\prime}):=\{f\in\mathcal{F_{\gamma}}(L):f(\cdot,\boldsymbol{c})\in\mathcal{F}^{\prime}\ \text{for all}\,\boldsymbol{c}\in[0,1]^{p}\}$. We prove a static-to-contextual regret conversion theorem that provides an upper bound for the contextual regret of the output algorithm as a function of the static regret of the input algorithm. Our construction of the output algorithm consists in discretizing the space of contexts into cells and running an algorithm $\pi$ separately within each cell. Similar constructions can be traced back to Hazan and Megiddo (2007), where the proposed method was not a meta-algorithm valid for any $\pi$ and $\mathcal{F}^{\prime}$ but a single contextual bandit algorithm with a specific input algorithm $\pi$ (projected gradient) and a specific class $\mathcal{F}^{\prime}$ (convex and Lipschitz functions). The analysis given in Hazan and Megiddo (2007) focuses on the $L$-regret and does not provide a control of the contextual regret. The subsequent work Cesa-Bianchi et al. (2017) adopts a similar discretization strategy, with similar limitations. It derives a bound on the $L$-regret when the input algorithm is chosen as Exp3 and the class $\mathcal{F}^{\prime}$ is the set of all 1-Lipschitz functions.
Another approach to static-to-contextual regret conversion is developed in Slivkins (2014). It relies on zooming with dyadic splits and is valid, again, for a specific input class $\mathcal{F}^{\prime}$, namely, a subset of Lipschitz functions, and a specific class of input algorithms $\pi$, namely, those achieving the optimal rate on this particular class. The class of functions $f(\cdot,\cdot)$ considered in Slivkins (2014) is a subclass of $\mathcal{G}_{1}(\mathcal{F}^{\prime})$ satisfying, along with the Lipschitz condition, the additional restrictive condition that for each $\boldsymbol{c}$, there exists $\boldsymbol{z}^{*}(\boldsymbol{c})=\arg\min_{\boldsymbol{z}\in\Theta}f(\boldsymbol{z},\boldsymbol{c})$ such that $\boldsymbol{c}\mapsto f(\boldsymbol{z}^{*}(\boldsymbol{c}),\boldsymbol{c})$ is Lipschitz. The algorithm of Slivkins (2014) requires exact knowledge of the Lipschitz constant, and the resulting bound on the contextual regret is exponential in the dimension of the context.
The contributions of the present paper can be summarized as follows.
  • (i)
    In Section 3, we propose a meta-algorithm that converts any input non-contextual bandit algorithm into an output contextual bandit algorithm. We prove a general conversion theorem that derives the contextual regret of the output algorithm from the static regret of the input algorithm (Theorem 2).
  • (ii)
    We apply our static-to-contextual conversion to handle the following two fundamental settings.
    • (ii-a)
      First, we consider the setting, where $f(\cdot,\cdot)$ is Lipschitz in both arguments, that is, $\gamma=1$ and $\mathcal{F}^{\prime}$ is the class of Lipschitz functions. In this setting, when the noises $\xi_{i}$ are zero, we show that our conversion approach leads to an output algorithm with a rate optimal contextual regret improving upon Slivkins (2014) in several respects. Unlike Slivkins (2014), we do not require the above restrictive additional condition on $f$ nor the knowledge of Lipschitz constant, and our bound scales polynomially in the dimension of the context.
    • (ii-b)
      Second, we consider the setting where $\mathcal{F}^{\prime}$ is the class of all convex functions and the noise in observations (2) is sub-Gaussian. In this case, by applying our conversion theorem together with the static regret bound from Fokkema et al. (2024) we prove that the contextual regret of the proposed output algorithm achieves the minimax optimal rate in $T$ up to a logarithmic factor.
  • (iii)
    In Section 4, we apply our conversion theorem in the setting where $\mathcal{F}^{\prime}$ is the class of strongly convex and smooth functions, and the noise is sub-Gaussian. We first construct a bandit algorithm, see Algorithm 2, which is an extension of the Bandit Convex Optimization (BCO) algorithm from Hazan and Levy (2014) to the problem with noisy queries. We prove a bound on its static regret. Then, combining it with the conversion theorem of Section 3, we derive a bound on the contextual regret of the proposed output algorithm.
  • (iv)
    In Section 5, we establish a minimax lower bound leading to several useful corollaries. First, it implies that the contextual regret bounds from items (ii-b) and (iii) are minimax optimal in $T$ up to logarithmic factors. Second, it shows that if $f(\boldsymbol{x},\cdot)$ lacks continuity, obtaining a sub-linear contextual regret may be impossible.
@@ 2 Related work
Contextual bandit problem has been introduced by Woodroofe (1979) in the framework of multi-armed bandits, that is, when the space of actions is finite. The literature on the contextual bandit problem with finite number of arms is now very rich, see, e.g., Goldenshluger and Zeevi (2009); Wang et al. (2005); Bastani and Bayati (2020); Perchet and Rigollet (2013); Rigollet and Zeevi (2010); Pandey et al. (2007); Langford and Zhang (2007); Auer et al. (2002); Kakade et al. (2008); Yang and Zhu (2002); Foster and Rakhlin (2023); Cella et al. (2023, 2022) and the references therein. In this literature, it is not uncommon to assume that the contexts are i.i.d. rather than adversarial.
There was also a considerable work on the problem of contextual bandits with continuous action space: Hazan and Megiddo (2007); Slivkins (2014); Kleinberg et al. (2008); Cesa-Bianchi et al. (2017); Chen and Gallego (2018); Li et al. (2019). These papers deal with adversarial contexts and, with the exception of Li et al. (2019), consider the noiseless setting ($\xi_{i}\equiv 0$). The analysis in Slivkins (2014) addresses the case where the context and the action vary in metric spaces with covering dimensions $p$ and $d$, respectively, under the assumption that $f(\cdot,\cdot)$ is Lipschitz in both arguments and satisfies an additional strong restriction (see the Introduction). The results in Slivkins (2014) provide a lower bound of the order $O(T^{\frac{p+d+1}{p+d+2}})$ on the contextual regret and an algorithm based on zooming techniques that attains this rate up to a logarithmic factor. Bounds for the $L$-regret for continuum contextual bandit algorithms are developed in Hazan and Megiddo (2007); Cesa-Bianchi et al. (2017). The paper Chen and Gallego (2018) deals with the setting where $f(\cdot,\cdot)$ is Lipschitz in both arguments, the actions belong to $[0,1]$, the function $f(\cdot,\boldsymbol{c})$ has a unique minimizer for any context $\boldsymbol{c}$ and exhibits approximately quadratic behavior around the minimizer. Under these assumptions, Chen and Gallego (2018) propose an algorithm with $O(T^{\frac{p+2}{p+4}})$ contextual regret. In Li et al. (2019), the authors addressed the problem of contextual bandits when the functions are strongly convex and smooth and derived a bound for the contextual regret that scales as $O(T^{\frac{p+1}{p+2}})$ up to logarithmic factors. To get this result, Li et al. (2019) additionally assumed that, for each $\boldsymbol{c}\in[0,1]^{p}$, the global minimizer $\operatorname*{arg\,min}_{\boldsymbol{x}\in\mathbb{R}^{d}}f(\boldsymbol{x},\boldsymbol{c})$ is an interior point of the constraint set $\Theta$. While this assumption might be hard to check in practice, it does simplify the analysis considerably. In the present paper, we drop this assumption and obtain in Section 4 an upper bound that improves upon the result of Li et al. (2019). Furthermore, in Section 5 we present a matching lower bound demonstrating that the algorithm that we propose is minimax optimal to within a logarithmic factor.
@@
In all the papers on continuum contextual bandits cited above, the contexts are adversarial, so the contextual regret is a variant of the dynamic regret as discussed after (3). The dynamic regret was also in the focus of studies in various settings of online convex optimization. For zero-order optimization (aka gradient-free optimization), the dynamic regret has been studied in Zinkevich (2003); Yang et al. (2016); Zhang et al. (2018); Chen and Giannakis (2018); Zhao et al. (2021). Among these papers, Yang et al. (2016); Chen and Giannakis (2018); Zhao et al. (2021) have addressed the bandit framework, in which the learner is not allowed to query function evaluations outside of the constraint set $\Theta$.
In the framework of adversarial multi-armed contextual bandits with finite number of arms, without imposing any regularity assumptions on the context such as our Assumption 1, Chen et al. (2019); Luo et al. (2018); Auer et al. (2019) achieved sub-linear regret by assuming that the functions $f_{t}$ either change infrequently or vary negligibly over time.
3 A general conversion theorem
In this section, we propose a meta-algorithm that to any input algorithm achieving a sub-linear static regret in adversarial bandit setting associates a contextual bandit algorithm with sub-linear contextual regret. We establish a conversion theorem (Theorem 2 below) that provides an upper bound for the contextual regret of the output algorithm as a function of the static regret of the input algorithm.
Here and in the sequel, unless other assumption is explicitly stated, we assume that the context $\boldsymbol{c}_{t}$ is drawn by an adversary who may be aware of all previous actions of the learner and past information received by the learner. We also assume throughout that for any $\boldsymbol{c}\in[0,1]^{p}$ the minimum $\min_{\boldsymbol{z}\in\Theta}f(\boldsymbol{z},\boldsymbol{c})$ is attained by some $\boldsymbol{z}^{*}(\boldsymbol{c})\in\Theta$.